# MegaScale-Infer: Serving Mixture-of-Experts at Scale with Disaggregated Expert Parallelism

### v2

@@ 5 High-Performance M2N Communication
M2N Sender. Figure 6 depicts the components of an M2N sender. In order to comply with the stream-oriented programming model, M2N senders launch low-utilization kernels that ensure the tensors to be transmitted are properly populated by preceding computation kernels. Once the tensors are ready, we utilize our in-house CPU communication library (denoted as Core Sender in the figure) which leverages RDMA write with immediate [10] and GPUDirect [12] for efficient transmission. This design eliminates complex GPU synchronization, GPU-to-CPU copies, and group initialization overhead, all of which contribute to significant latency issues, especially for relatively small tensor sizes, as demonstrated in Figures 10 and 11.
M2N Receiver. Figure 7 illustrates the components of an M2N Receiver. Since communication between M2N senders and receivers leverages pre-registered RDMA buffers, the receiver’s primary task is relatively straightforward—ensuring that tensors have been successfully transmitted from the corresponding senders. This can be efficiently achieved by polling the relevant completion queue [9]. Additionally, for ease of integration and testing, we provide a custom copy kernel that optionally copies the received tensors from the pre-registered buffers to user-defined output tensors arranged in a predefined layout.
@@ 6 Implementation
High-performance M2N communication library. We built our communication library as a Pytorch extension [14] in around 4900 and 5000 lines of C/C++ and Python code, respectively. Our library is supported by technologies such as GPUDirect and GDRCopy [11] (utilized by M2N receivers when custom copy kernels are used). We also carefully design network monitoring tools to delve into network and traffic-related optimizations.
@@ 7.1 Experimental Setup
Baselines. We compare MegaScale-Infer with two state-of-the-art serving systems: vLLM [43] and TensorRT-LLM [50]. Both systems support popular techniques for LLM serving, including FlashAttention [26], PagedAttention [43], and continuous batching [71]. They primarily rely on tensor parallelism for distributed LLM serving, with TensorRT-LLM additionally supporting expert parallelism for expert layers. For larger models requiring inter-node communication, both systems also support pipeline parallelism. Due to GPU memory limits and large model sizes, serving Mixtral 8x22B and DBRX with these baselines requires a minimum of 8 GPUs, while the scaled-MoE necessitates multi-node deployment.
@@ 7.3 Performance of M2N Communication
Figure 10 illustrates how the latency and throughput of M2N vary with different data sizes. In this experiment, we set the number of senders and receivers to 8. Overall, MegaScale-Infer achieves lower latency and higher throughput than NCCL across all data sizes. NCCL incurs substantial overhead for small data sizes due to additional data copies and group operations. To mitigate these overheads, we design and implement highly optimized communication kernels with GPUDirect, resulting in up to an 80.8% reduction in median latency, as shown in Figure 10(a). Additionally, NCCL suffers from high tail latency due to its inherent instability. We address this issue by eliminating GPU synchronization and group initialization. As depicted in Figure 10(b), we achieve up to 96.2% reduction in P99 latency compared to NCCL. These optimizations enable MegaScale-Infer to improve throughput by up to $\times9$. For commonly used data sizes in model serving, such as 256KB, MegaScale-Infer achieves improvements of 68.2% in median latency, 92.9% in tail latency, and a 4.2×4.2$\times$4.2 × increase in throughput.

### v3

@@ 5 High-Performance M2N Communication
M2N Sender. Figure 6 depicts the components of an M2N sender. In order to comply with the stream-oriented programming model, ① M2N senders launch low-utilization kernels which ensures that the ② pre-registered tensor to be transmitted is properly populated by preceding computation kernels. ③ Once the tensor is ready, ④ our in-house CPU communication library (denoted as Core Sender in the figure) transmits the tensor efficiently using RDMA write with immediate [11]. Finally, to guarantee proper utilization of the registered tensor, ⑤ the sender polls completions from the corresponding completion queue  [10], confirming that data has been written to the remote buffers. This design eliminates complex GPU synchronization, GPU-to-CPU copies, and group initialization overhead, all of which contribute to significant latency issues, especially for relatively small tensor sizes, as demonstrated in Figures 10 and 11.
M2N Receiver. Figure 7 illustrates the components of an M2N Receiver. Since communication between M2N senders and receivers leverages ⑥ pre-registered RDMA buffers, the receiver’s primary tasks are relatively straightforward. First, it must verify that tensors have been successfully transmitted from the corresponding senders, which is efficiently achieved by ⑦ polling the relevant completion queue. Second, to ensure data consistency at the GPU level, our in-house CPU communication library (denoted as Core Receiver in the figure) leverages GDRCopy [12] and performs a ⑧ flush operation [1]. Finally, the receiver ensures that the subsequent kernel in the compute stream waits until the data has been correctly placed in the destination buffer, which is achieved by ⑨ updating a shared flag. This simpler design eliminates the need for GPU-to-GPU copies and effectively reduces the GPU utilization overhead of receivers.
@@ 6 Implementation
High-performance M2N communication library. We built our communication library as a Pytorch extension [15] in around 4900 and 5000 lines of C/C++ and Python code, respectively. As discussed earlier, our library is supported by technologies such as GPUDirect [13] and GDRCopy. We also carefully design network monitoring tools to delve into network and traffic-related optimizations.
@@ 7.1 Experimental Setup
Baselines. We compare MegaScale-Infer with two state-of-the-art serving systems: vLLM [44] (version 0.6.6.post1) and TensorRT-LLM [51] (version 0.15.0). Both systems support popular techniques for LLM serving, including FlashAttention [27], PagedAttention [44], and continuous batching [72]. They primarily rely on tensor parallelism for distributed LLM serving, with TensorRT-LLM additionally supporting expert parallelism for expert layers. For larger models requiring inter-node communication, both systems also support pipeline parallelism. Due to GPU memory limits and large model sizes, serving Mixtral 8x22B and DBRX with these baselines requires a minimum of 8 GPUs, while the scaled-MoE necessitates multi-node deployment.
@@ 7.3 Performance of M2N Communication
Figure 10 illustrates how the latency and throughput of M2N vary with different data sizes. In this experiment, we set the number of senders and receivers to 8. Overall, MegaScale-Infer achieves lower latency and higher throughput than NCCL across all data sizes. NCCL incurs substantial overhead for small data sizes due to additional data copies and group operations. To mitigate these overheads, we design and implement a highly optimized communication library, resulting in up to an 80.8% reduction in median latency, as shown in Figure 10(a). Additionally, NCCL suffers from high tail latency due to its inherent instability. We address this issue by eliminating GPU synchronization and group initialization. As depicted in Figure 10(b), we achieve up to 96.2% reduction in P99 latency compared to NCCL. These optimizations enable MegaScale-Infer to improve throughput by up to $\times9$. For commonly used data sizes in model serving, such as 256KB, MegaScale-Infer achieves improvements of 68.2% in median latency, 92.9% in tail latency, and a 4.2×4.2$\times$4.2 × increase in throughput.